I'm performing new analysis on the data taken in February and in May during the shifts in which we changed the injected squeezing level into the ITF by changing the green pump power. The descripiton of the activity performed during these shifts can be found in the entries #44846 and #45918.
If we plot the spectrum of hrec with different pump power level (i.e. different level of injected squeezing) we can see that in February (fig 1) we measured anti-squeezing from 35 Hz whereas in May (fig 2) only from 55 Hz.
I remember that in February we had the old photodiodes whereas in May the HQE photodiodes.
In February we could see anti-squeezing (so also squeezing) from lower frequencies, this could be one of the limiting factors of why in the last months the sensitivity of the interferometer increases less injecting squeezing.
A question: less anti-squeezing can be due to higher losses in the squeezing injection, thus this effect can be explained with higher noise at low frequencies caused by the photodioed electronic flicker noise?